When aluminum reacts with hydrochloric acid, HCl, both aluminum chloride and hydrogen gas are produced. If 108 grams of aluminum is reacted, what volume of gas may be collected at STP?


1
Expert's answer
2021-04-11T23:49:03-0400

Solution:

Balanced chemical equation:

2Al + 6HCl → 2AlCl3 + 3H2

According to the equation above: n(Al)/2 = n(H2)/3


The molar mass of Al is 27 g/mol.

Hence,

Moles of Al = (108 g Al) × (1 mol Al / 27 g Al) = 4 mol Al

n(Al) = 4 mol


n(H2) = 3 × n(Al) / 2 = (3 × 4 mol) / 2 = 6 mol


At STP, one mole of any gas occupies a volume of 22.4 L.

Thus 6 mol of H2 occupies:

(6 mol × 22.4 L) / 1 mol = 134.4 L of H2


Answer: 134.4 L of H2 may be collected at STP.